One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
a)5
b)7
c)17
d)34
I am usually good with these but I am having a hard time with this one, and need some help please. Thanks.

Answer by Earlsdon(6103)

Let the two numbers be x and y, with the smaller number being x and the larger number being y.
You can write:
y = x+12 Because y is the larger and it's 12 more than the other number.
x + 2y = 39 Substituting the y from above:
x + 2(x+12) = 39 Simplifying:
3x+24 = 39
3x = 15
x = 5 This is the smaller number.
y = x+12
y = 5+12
y = 17 This is the larger number.